3) Two 1.0 kg masses are 1.0 m apart (center to center) on a frictionless table. Each has a +10uC of charge. a) What is the magnitude of the electric force on one of the masses? b) What is the initial acceleration of one of the masses if it is release<d and allowed to move?
Answer #1

a) electric force = kq1q2 / r^2

where q1,q2 are the charges

r = separation

F = 9*10^9*(10*10^-6)^2 / 1^2

= 0.9 N

b) accleration = Force / mass

= 0.9 / 1 = 0.9 m/s2